### v2

# Evaluating LLMs’ Mathematical Reasoning in Financial Document Question Answering

Abstract
Large Language Models (LLMs), excel in natural language understanding, but their capability for complex mathematical reasoning with an amalgamation of structured tables and unstructured text is uncertain. This study explores LLMs’ mathematical reasoning on four financial tabular question-answering datasets: TATQA, FinQA, ConvFinQA, and Multihiertt. Through extensive experiments with various models and prompting techniques, we assess how LLMs adapt to complex tables and mathematical tasks. We focus on sensitivity to table complexity and performance variations with an increasing number of arithmetic reasoning steps. The results provide insights into LLMs’ capabilities and limitations in handling complex mathematical scenarios for semi-structured tables. Ultimately, we introduce a novel prompting technique tailored to semi-structured documents, matching or outperforming other baselines in performance while providing a nuanced understanding of LLMs abilities for such a task.
1 Introduction
In the constantly evolving realm of artificial intelligence, Large Language Models (LLMs) have risen as cutting-edge tools for natural language understanding. They excel in a wide array of NLP tasks, including machine translation (MT), text summarization, question answering, and code generation. One specific area where LLMs’ mathematical reasoning abilities come under scrutiny is the domain of numerical reasoning tasks. Past research has delved into the potential of language models for mathematical reasoning tasks, as seen in studies such as Amini et al. (2019); Upadhyay and Chang (2017); Patel et al. (2021); Cobbe et al. (2021). These investigations provide a means to evaluate the performance of language models when it comes to solving mathematical problems, ranging from straightforward math word problems to more complex ones.
Figure 1: An example of a QA pair on a financial table.
However, the problem becomes significantly more challenging when we encounter a hybrid of structured and unstructured data, as exemplified by semi-structured tables. Such tables are common in documents like invoices, health records, and financial reports in corporate settings. In the financial domain, these tables present numerical data in a structured format, including income statements, balance sheets, cash flow statements, shareholder equity data, and annual reports. The majority of NLP models are primarily trained to handle unstructured textual data, which limits their ability to reason over semi-structured data, such as tables, or more intricate hybrids of tables and text, as seen in Chen et al. (2020b); Aly et al. (2021); Chen et al. (2020a, 2021a). Tables, especially those with intricate relationships and dependencies, often necessitate multi-hop reasoning, connecting information across multiple steps.
NLP models may encounter difficulties in performing such multi-step reasoning, particularly when dealing with complex mathematical operations involving tables, as highlighted in Li et al. (2022). Previous research Chen (2023), exemplified these issues and demonstrated LLM capacity to process and reason with semi-structured tables. However, these studies are somewhat constrained and don’t explicitly explore the models’ mathematical reasoning abilities. This is particularly evident in data/tasks that involve a substantial number of arithmetic reasoning steps, operate on extreme orders of magnitude, or deal with intricate tables where extracting the relevant information for a query becomes challenging.
Moreover, when handling domain-specific documents, such as those in finance, a language model must not only have the necessary domain knowledge to craft the right approach for task-solving but also the capability to manipulate structured data, such as tables. Therefore, in this study, we aim to fill this gap by providing both qualitative and quantitative analyses of LLM’s ability to reason on mathematical content on four finance datasets. Our goal is to illustrate how model performance varies as the numerical complexity of the underlying data and the intricacy of mathematical reasoning steps required to solve a query increase. Our contributions are as follows:
1. We conduct a comprehensive evaluation of state-of-the-art Large Language Models (LLMs) in tabular question answering, with a specific focus on mathematical reasoning tasks, using public financial tabular datasets to establish a robust performance benchmark.
2. Our analysis is thorough and multifaceted, encompassing both qualitative and quantitative aspects. We aim to provide nuanced insights into the strengths and limitations of LLMs in tabular question answering, especially in scenarios involving mathematical reasoning.
3. Building upon qualitative analysis, we introduce a novel and improved prompting method called EEDP (Elicit $\longrightarrow$ Extract $\longrightarrow$ Decompose $\longrightarrow$ Predict). This innovative approach yields comparable or enhanced performance when compared to existing prompt engineering methods. This method of explicitly prompting the model to break it’s response into sections provides a more explainable and nuanced understanding of their weaknesses in performing such tasks.
2 Related Works
2.1 LLMs on Mathematical Reasoning
Pre-trained Language Models (PLMs) excel in NLP tasks Devlin et al. (2019); Zhuang et al. (2021) by leveraging extensive textual corpora to acquire world knowledge Guu et al. (2020). Expanding PLMs for math-related tasks has been challenging due to their non-specific training. Recent attempts include MWP-BERT and Minerva Liang et al. (2022); Lewkowycz et al. (2022), but curating high-quality math data remains difficult. To bridge the gap, researchers fine-tune PLMs for specific math tasks. Notable works, like Bhaskara, Self-sampler, Aristo, FinQANet, TAGOP, MT2Net, and others Mishra et al. (2022); Ni et al. (2022); Clark et al. (2021); Chen et al. (2021b); Zhu et al. (2021); Zhao et al. (2022); Cao and Xiao (2022); Welleck et al. (2022), employ PLMs such as GPT-Neo and RoBERTa for math problem-solving.
2.2 Tabular Question Answering
Handling diverse input formats in question answering, including structured tables and visual data, poses challenges for language models. HybridQA Chen et al. (2020b) introduces questions requiring reasoning over tables and text. MultimodalQA Talmor et al. (2021) adds visual inference. Our focus is on multi-hop question answering over tables and text. TAPAS Herzig et al. (2020) tackles table-based questions without logical forms, while Tapex Liu et al. (2022) empowers generative models with table reasoning. Models like FinQANet, TagOP, and MT2Net Chen et al. (2021b); Zhu et al. (2021); Zhao et al. (2022) consist of a retriever module to first retrieve the supporting facts from the input financial report followed by a reasoning module to perform reasoning over the retrieved facts and get the final answer. Both Retrieving the right evidences from financial reports and reasoning over the retrieved facts require in-domain knowledge. For instance, questions may require understanding the financial implications of transactions, identifying trends in revenue or interpreting complex statements regarding investments. This necessitates a reliance on an external knowledge base to overcome the challenges faced while dealing with domain-specific multi-hop question answering.
2.3 Prompt Engineering
In-context Learning (ICL) equips Language Models (LLMs) with task examples and queries, enabling them to perform target tasks without updating model parameters (Brown et al., 2020; OpenAI, 2023). They excel in mathematical reasoning with few-shot prompts but struggle with more complex tasks. Methods like chain-of-thoughts (CoT) (Wei et al., 2022) have improved LLM performance by guiding them through intermediate reasoning steps. Enhancing multi-step reasoning involves two main approaches: improving in-context examples and obtaining better reasoning steps. Some focus on more stable in-context example selection (Fu et al., 2023; Rubin et al., 2022; Lu et al., 2023). Others adopt a modular approach, using off-the-shelf tools (Schick et al., 2023), program of thoughts (PoT) (Chen et al., 2022a), or decomposition strategies (Zhou et al., 2023; Dua et al., 2022; Khot et al., 2023). Our task requires complex multi-step reasoning across diverse information sources. LLMs, as demonstrated by (Chen, 2023), can reason over structured tables without explicit encoding. They also serve as versatile decomposers, breaking down extensive evidence and complex questions (Ye et al., 2023b).
3 Dataset Annotations
In our evaluations, we considered four tabular datasets: FinQA Chen et al. (2021b), TATQA Zhu et al. (2021), ConvFinQA Chen et al. (2022b), and Multihiertt Zhao et al. (2022). These datasets involve questions requiring numerical reasoning over semi-structured data. We excluded simple span selection questions from TATQA. Each dataset provides pre-annotated executable programs for deriving ground-truth answers. We meticulously annotate the datasets considered for our study with the following metadata:
1. Number of Reasoning Steps:
Including the count of arithmetic operations in questions is crucial. More operations reflect increased complexity in reasoning, and their interdependence offers insights into the models’ proficiency. This annotation, applied across all four datasets, reveals their ability in handling intricate arithmetic tasks.
Refer to Figure 5 in Appendix (A.2), answering the question require multiple reasoning steps.
2. Question Categorization:
In numerical reasoning, grasping the evolution from fundamental arithmetic to advanced operations is crucial, marking a shift in cognitive complexity. As questions advance, they typically involve more intricate combinations of operations and linguistic nuances. Our research seeks to identify both the capabilities and limitations of LLMs in understanding these concepts.
We establish 12 mathematical concept categories (Table 1) with corresponding definitions, annotating each question. The dataset coverage across these categories is shown in Figure 6 in Appendix (A.2). Notably, categories like Division and Ratio share similarities but differ in focus: Division involves the division operator, while Ratio encompasses ratios, fractions, and inverse problems. Change in Ratio questions add complexity with quantity changes requiring subtraction. Additionally, we omit NEED-IN-DOMAIN-INFO due to domain-specific knowledge focus and TIME questions due to limited sample size.
Concepts Definition Sum Questions that require only the knowledge of addition. Difference Questions that require only the knowledge of subtraction. Product Questions that require only the knowledge of multiplication. Division Questions that require only the knowledge of division. Ratio Questions that require knowing fractional forms, e.g., percentages, ratios. Change Ratio Questions involving the difference between two fractional forms, e.g., percentage changes, difference in ratios. Range Questions requiring knowledge of the minimum and maximum of data observations. Compare Questions necessitating a comparison between mathematical quantities (e.g., greater than, less than). Average Questions needing knowledge of the average, used to calculate the central tendency of a group of data points. In-Domain-Info Questions that require implicit knowledge to understand domain-specific mathematical formulations (e.g., return on investment (RoI), cost of goods sold (COGS), amortization rate, etc.). Time Questions explicitly involving mathematical operators for time-spans not in the table or context. Counting Questions requiring the counting of elements in a set or group of data points.
Table 1: Mathematical Concept Categories and Definitions for Studying Concept Comprehension Abilities of LLMs
3. Number of Rows:
Evaluating performance with larger supporting tables is crucial. Larger tables complicate multi-hop reasoning tasks by increasing the amount of information, making it harder to identify relevant evidence. We prioritize these annotations for datasets like FinQA and Multihiertt, where questions mainly use tables as supporting evidence. Therefore, these annotations are confined to these datasets. In Multihiertt, when multiple tables support evidence, we consider the one with the highest row count. The dataset coverage across the number of rows for Multihiertt and FinQA is shown in Figure 7 in Appendix (A.2).
4. Depth of Hierarchy:
In hierarchical tables, such as those in Multihiertt, evaluating model performance concerning the growing depth of hierarchy in cells with critical information becomes paramount. To tackle this, we annotate each example in Multihiertt with the hierarchy depth of cells containing crucial information, considering the maximum hierarchy value for multiple value extraction. Our approach to estimating hierarchy depth is illustrated in Figure 16, and Figure 8 in Appendix (A.2) displays the sample distribution that require the extraction of gold evidence from various hierarchy depths.
5. Proportion of Empty Cells:
Interpreting a table becomes challenging as the number of empty cells increases. Empty cells indicate missing or undefined information, leading to potential gaps in understanding.
Assessing empty cell proportions is crucial to quantify data ambiguity. More empty cells suggest higher uncertainty, which can hinder models’ ability to derive meaningful insights and impact reasoning accuracy. In Multihiertt, where tables are hierarchical in nature and empty cells occur quite frequentlt, we annotate examples with the percentage of empty cells, contributing to our understanding of data ambiguity. For the distribution of empty cells, refer to Figure 9 in Appendix (A.2).
4 Model Performance
Models:
In this study, we choose to experiment with LLMs like GPT-3.5-Turbo, GPT-4, PaLM-540B, Mistral-7B-Instruct [^1] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, Llama-2-13B [^2] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf and MAmmoTH-13B [^3] https://huggingface.co/TIGER-Lab/MAmmoTH-13B.
Baseline Prompting Strategies:
Prompt Engineering involves designing inputs for LLMs to produce optimal results. Typically, for an Instruction-tuned LLM, it’s assumed that we give the model a task-specific instruction $\mathcal{I}$ accompanied with a few (usually $k\in\{2,4\}$) demonstrations $\mathcal{D_{T}}$ of a task T. We experiment with the following prompting techniques as baselines.
  • Direct: In this setup, we explicitly instruct the models to abstain from providing explanations and just return the final answer to the posed question. For this scenario, $\mathcal{D_{T}}$ contains $\{(p_{i},q_{i},a_{i})\}_{i=1}^{k}$ where p is the premise which includes the table and the associated text, q is the question and a is the ground-truth answer.
  • CoT: Moving forward, we experiment with the chain-of-thoughts prompting strategy where we instruct the model to output the explanation to the answer derived by it. Here, our $\mathcal{D_{T}}$ contains $\{(p_{i},q_{i},e_{i})\}_{i=1}^{k}$ where p is the premise which includes the table and the associated text, q is the question and e is the explanation of the answer.
  • PoT: In this case, the expected response is a code derivation of the final answer. Here, our $\mathcal{D_{T}}$ contains $\{(p_{i},q_{i},c_{i})\}_{i=1}^{k}$ where p is the premise which includes the table and the associated text, q is the question and c is the code-derivation of the answer.
  • Decomposers: Ye et al. (2023a) proposed to address the challenge of handling large tables by decomposing them into more manageable subtables. Similarly, we handle complex questions by breaking them down into simpler subquestions. In their work, they show that their method works well on SQL tables where the table can easily be shortened to remove distractive information without the loss of any supporting evidence. Next, they parse the question to decompose it into simpler, more manageable subquestions. Finally, the model is expected to answer each of the subquestions and compose those to arrive at the final answer. In this case, our demonstration set $\mathcal{D_{T}}$ contains $\{(p_{i}^{{}^{\prime}},\langle q_{1},q_{2},...,q_{n}\rangle,a_{i})\}_{i=1}^{k}$ where $p^{{}^{\prime}}$ is the premise obtained by the removal of information irrelevant to the question from the original premise p and $\langle q_{1},q_{2},...,q_{n}\rangle$ are the subquestions whose answer lead to the final answer.
EEDP (Ours):
We propose a novel prompting strategy: Elicit $\longrightarrow$ Extract $\longrightarrow$ Decompose $\longrightarrow$ Predict.
  1. 1.
    Elicit: We prompt the model explicitly to first elicit relevant domain knowledge for answering a given query.
  2. 2.
    Extract: Conditioned on the table, question and the elicited domain knowledge, the model extracts supporting evidences to answer a given question.
  3. 3.
    Decompose: We instruct the LLM to break a complex mathematical reasoning task into multiple atomic operations and compose the operations to arrive at the final answer.
  4. 4.
    Predict: The model finally returns the derived answer in the above steps,
Comparison with other state-of-the-art methods
  1. 1.
    PoT vs EEDP: PoT enhances LLM inference with the use of variable names for the supporting values extracted from the premise and prompts the LLM to express their thought process in the form of programs. The model output is a program which is executed externally to derive the final answer. EEDP proposes to decompose a complex reasoning task into simple atomic steps whose solutions can be composed to give the final answer. In PoT, they don’t make the language model do the computation while in our case the language model not only outputs the reason but also computes the final answer. This distinction implies that PoT may have an inherent advantage over EEDP.
  2. 2.
    Decomposers vs EEDP: The prompting strategy proposed by Ye et al. (2023b) was originally designed for quering SQL tables, they use LLMs to break down evidence and questions for SQL interpreters. In contrast, our approach addresses more complex scenarios involving both tables and text, requiring advanced reasoning skills beyond the capability of standard SQL interpreters. Pruning a non-SQL table using this method can lead to significant information loss from the premise which can be a potential ingredient required to derive the final answer. Additionally, this is an expensive method as it requires 3X API calls as opposed to other prompting methods. Moreover, EEDP is a unified prompting strategy which integrates multiple solver elements into a single unified prompt for elicitation, extraction, decomposition and prediction.
Table 2. The reason behind this was statistically found to be inaccurate formation of subtables from the main table leading to information loss as described in the previous paragraph. The performance of EEDP either surpasses or matches very closely with that of PoT. The number of shots were adjusted depending upon the context length of the model. We can see that MammoTH-13B which is finetuned on the MathInstruct dataset Yue et al. (2023) containing Instruction-Response pairs where the responses are a hybrid of CoT and PoT rationales fails to do well on this task with EEDP methodology. We argue that this is due to two potential reasons:
  1. 1.
    Reduction of the number of shots to adjust the context length as the EEDP response is longer than that of the other methods.
  2. 2.
    Finetuning may contribute to suboptimal performance due to its alignment with a particular style and format of responses, potentially limiting the model’s adaptability and generalization to diverse contexts.
Fig. 17 shows a template for EEDP strategy. Table 2 shows a comparison in performance between different prompting strategies.
Model Dataset Direct CoT PoT EEDP Decomposers
TATQA 55.81 86.91 89.99 88.67 47.46
GPT-4 FinQA 65.12 72.38 75.26 76.05 44.93
ConvFinQA 63.10 71.19 78.81 77.91 18.76
Multihiertt 41.35 55.13 67.23 70.32 36.86
TATQA 31.38 77.57 82.11 79.73 28.53
GPT-3.5-Turbo FinQA 40.47 59.18 68.97 61.88 32.33
ConvFinQA 37.62 48.33 61.19 61.75 10.50
Multihiertt 25.88 42.33 52.18 49.65 20.61
TATQA 44.66 62.93 61.60 81.51 57.94
PaLM 2-540B FinQA 30.33 34.79 30.41 61.95 46.38
ConvFinQA 20.19 38.00 40.14 63.42 22.32
Multihiertt 14.20 20.67 36.52 37.97 20.19
TATQA 3.36 35.95 34.16 40.95 25.93
Llama 2-13B FinQA 1.80 25.34 12.97 30.47 11.91
ConvFinQA 3.80 29.45 29.92 39.42 10.35
Multihiertt 1.54 30.66 18.12 24.15 16.86
TATQA 19.11 56.25 10.02 4.37 22.89
MammoTH-13B FinQA 22.83 35.32 15.86 35.05 17.65
ConvFinQA 21.61 46.08 8.78 32.77 7.83
Multihiertt 10.12 18.56 6.57 18.36 11.87
TATQA 10.92 59.14 16.53 56.06 7.24
Mistral-7B FinQA 26.11 34.23 10.56 34.86 12.34
ConvFinQA 12.35 48.45 14.48 36.57 11.16
Multihiertt 14.909 22.92 14.94 10.97 11.63
Table 2: Comparison of performance of different models tested against a variety of prompting strategies
RQ: Where do LLMs falter while performing hybrid question answering involving arithmetic operations?
We categorize errors and identify their origins through a detailed analysis of model responses.
1. Incorrect Extraction:
This category encompasses errors where the model faces difficulties in accurately identifying and extracting the pertinent information necessary for effective problem-solving. These errors point to challenges in retrieving precise information. These errors can further be subdivided into two categories
  • Missing/Incomplete Evidences: The model fails to extract all the necessary evidences which serve as ingredients to derive the final answer.
  • Wrong Evidences: The model extracts wrong values for variables as supporting evidences from the premise.
2. Incorrect Reasoning:
Errors in reasoning occur when the model struggles to formulate an appropriate and contextually relevant approach to tackle a given problem. Possible reasons include a lack of domain knowledge or an inaccurate interpretation of the posed question. Consequently, this error type can arise from two sources.
  • Deficit in Domain Knowledge: These errors occur when the model attempts to derive an answer to the posed question using a wrong formula for domain-specific measures, for eg. COGS, ROI etc.
  • Question Misinterpretation: These errors occur when the model interprets the question differently and provides responses that are not aligned with the intended query. Overall, the model’s outputs lack grounding in the original question posed to it.
3. Incorrect Calculation:
This variety of errors include those where the model commits mistakes due to calculation mistakes. This can be of two types.
  • Incorrect Instantiation: These include cases if the model extracts the right evidences, uses the right derivation formula but instantiates the variables incorrectly with the values resulting in an incorrect answer.
  • Precision Error: Language models employ mathematical algorithms for arithmetic operations, but their results may not always be perfectly accurate due to insufficient data pattern coverage or introduced biases during training. Consequently, they can sometimes generate outputs with slight inaccuracies or deviations from correct results. We show a detailed analysis in A.1.
The above categorization provides a nuanced understanding of the diverse challenges and shortcomings exhibited in different facets of mathematical reasoning. We observe that in a lot of cases, the error propagates because of a deficiency in domain knowledge. It is critical for both evidence extraction and reasoning. Despite possessing general domain knowledge owing to the massive amount of data these models have been pre-trained upon, these models may require explicit prompts to elicit the specific domain knowledge needed for a particular question. Furthermore, errors can arise due to the models’ limited proficiency in multi-step reasoning, especially in tackling complex questions involving multiple arithmetic operations in a sequence.
We give a quantitative measure of each type of errors for each of the 4 datasets we consider for our study in Table 3. We also provide examples corresponding to each error category in figures 10, 11, 12, 13, 14 and 15. Statistically, we find that reasoning errors contribute a significant chunk to the total number of errors. In case of complex hierarchical tables like that in Multihiertt, the model is found to struggle with extracting the right supporting evidences from the premise for a given question. Calculation errors can be taken care of if a third-party calculation tool is chained to the language model.
Dataset Error Type Perc. (%)
FinQA Extraction.C1 10.38
Extraction.C2 25
Reasoning.C1 25
Reasoning.C2 15.57
Calculation 24.06
ConvFinQA Extraction.C1 8.45
Extraction.C2 14.08
Reasoning.C1 19.72
Reasoning.C2 36.62
Calculation 21.13
TATQA Extraction.C1 13.79
Extraction.C2 31.03
Reasoning.C1 22.41
Reasoning.C2 5.17
Calculation 27.59
Multihiertt Extraction.C1 20.5
Extraction.C2 31.5
Reasoning.C1 15.5
Reasoning.C2 12
Calculation 20.5
Table 3: Error Analysis on Various Datasets. In this table, Extraction.C1: Missing Evidences, Extraction.C2: Wrong Evidences, Reasoning.C1: Insufficient Domain Knowledge, Reasoning.C2: Question Misinterpretation, Calculation: Instantiation and Precision errors
5 Analysis on Reasoning Annotations
We analyse model performance on the basis of fine-grained annotations as described in the section 3.
1. Performance vs Number of Reasoning Steps
[…]
Figure 2: A comparison showcasing the performance trends across various datasets with the increasing number of reasoning steps. The analysis contrasts the effectiveness of EEDP (our method) against PoT in addressing complex reasoning.
We investigate model performance with increasing mathematical reasoning steps, as shown in Figure 2. This analysis provides insights into models’ ability to handle varying task complexities. As expected, performance decreases with more reasoning steps, indicating LLMs’ challenges in retrieving information and reasoning as complexity grows.
Anomalies are observed in ConvFinQA, where accuracy improves with two reasoning steps, potentially due to questions referring to prior conversation turns. Anomalies like these warrant further investigation beyond this study’s scope.
2. Performance across Question Types
We explore the performance trend across distinct question categories, as defined in Table 1, to evaluate the models’ familiarity with various mathematical and financial concepts. Figure 3 demonstrates that EEDP exhibits a performance competitive to PoT for simpler datasets like FinQA, ConvFinQA and TATQA but has the potential to significantly improve the performance in case of a complex dataset like Multihiertt.
[…]
Figure 3: A comparison showcasing the performance trends observed in various datasets across different question types. The analysis contrasts the effectiveness of EEDP (our method) against Few-Shot PoT (PoT). Best viewed in color.
6 Conclusion
In conclusion, our study delved into LLMs’ mathematical reasoning in complex financial scenarios, assessing their adaptability to mixed structured tables and unstructured text. Through rigorous experimentation, we uncovered insights into their performance and limitations, presenting a tailored prompting technique that outperformed other baseline methods. Our findings advance understanding of LLMs’ abilities in tackling intricate mathematical tasks within semi-structured documents, suggesting directions for future research.
7 Limitations
The scope of this work is limited by the following challenges:
  1. 1.
    There are not many datasets dealing with numerical reasoning over semi-structured data apart from the ones from financial domain. As a future work, it would be interesting to similar analysis across various domains such as e-commerce, healthcare and scientific tables from research papers.
  2. 2.
    It is essential to acknowledge that there may be multiple factors influencing the performance of large language models while dealing with numerical reasoning over semi-structured data. In our analysis, we have focused on specific factors and trends, but it is important to recognize that the overall performance is affected by a multitude of variables. Marginalizing (observing the trend along one while keeping the rest as constants) or isolating a single factor is challenging and cannot be done with real-world data. Additionally, future investigations may benefit from simulating controlled scenarios on synthetic datasets to gain deeper insights into the impact of individual factors on model performance.
  3. 3.
    We emphasize our analysis on prominent models such as GPT-4, GPT-3.5-TURBO, and PaLM 2-540B due to their substantial size and capabilities. Notably, other open-sourced large language models like LLaMA 2-13B, MAmmoTH-13B and Mistral-7B-Instruct did not exhibit satisfactory performance in numerical reasoning over semi-structured data. This accentuates the need for exploring computationally feasible and cheaper models that can deliver remarkable performance in tasks involving numerical reasoning over heterogeneous source of information.
  4. 4.
    As a future work, it would be intriguing to integrate special position encodings (similar to those utilized in TAPAS) into state-of-the-art language models for every cell in the table. This integration would facilitate the extraction of information from the table, considering its structure more effectively.
  5. 5.
    As a future direction, models specifically tuned to answer to human queries over complex documents in multiple conversational turns (like that in ConvFinQA) should also be considered as it’s a challenge for language model’s to backtrack to their decisions which were made previously in the conversation.
8 Ethics Statement
We, the authors of this work, affirm that our work complies with the highest ethical standards in research and publication. In conducting this research, we have considered and addressed various ethical considerations to ensure the responsible and fair use of computational linguistics methodologies. We provide detailed information to facilitate the reproducibility of our results. This includes sharing code, datasets (in our case, we deal with publicly available datasets and comply to the ethical standards mentioned by the authors of the respective works.), and other relevant resources to enable the research community to validate and build upon our work. The claims in the paper match the experimentation results, however, with black-box large language models, a certain degree of stochasticity is expected which we attempt to minimize by keeping a fixed temperature. We describe in the fullest details the annotations, dataset splits, models used and prompting methods tried, ensuring reproducibility of our work.
Acknowledgements
Appendix A Appendix
A.1 How proficient are LLMs in performing simple arithmetic operations?
To assess the effectiveness of Large Language Models in handling fundamental arithmetic tasks (addition (+), subtraction (-), multiplication (*), and division (/)) across operands of varying magnitudes, we generate a set of 2600 synthetic arithmetic expressions using GPT-4. This set includes 650 problems for each arithmetic operation. Within each operation category, we categorize tasks into groups based on a parameter denoted as $\tau$:
[…]
where, $\arg\max$ selects the operand op with the greater absolute value of the order of magnitude, and OOM represents the order of magnitude.
This approach is motivated by cognitive challenges commonly faced by humans, as they often encounter difficulties with both high and low orders of magnitude. Essentially, captures the order of magnitude of the operand with the larger absolute value among the two. For each arithmetic operation, we establish groups with $\tau$, ranging from -6 to 6. Within each group, there are 50 arithmetic expressions. This systematic grouping provides a comprehensive assessment across various difficulty levels based on operand magnitudes.
Figure 4 shows that in case of relatively simpler arithmetic operations like addition and subtraction, the effect of order of magnitude of the operands is less profound as compared to that in case of harder arithmetic operations like multiplication and division. We observe the trend in the performance accuracy with the growing and diminishing orders of magnitude. We also observe bigger model such as GPT-4, GPT-3.5-TURBO and PaLM 2-540B perform much better on simpler addition, subtraction task in comparison to multiplication, division task.
Figure 4: Accuracy of different arithmetic operations across different orders of magnitude
A.2 Dataset Coverage across Reasoning Annotations
Figure 5: Sample distribution of examples in numerical reasoning on tabular datasets partitioned by the number of reasoning steps involved. Clockwise (from top-left) : FinQA, ConvFinQA, TATQA, Multihiertt.
Figure 6: Sample distribution of Numerical & Tabular Reasoning Datasets partitioned by Question Concept Category types. Clockwise (from top-left): FinQA, ConvFinQA, Multihiertt, TATQA
Figure 7: Sample distribution of Multihiertt & FinQA datasets partitioned by number of rows in the supporting table.
Figure 8: Sample distribution of Multihiertt Dataset partitioned by depth of hierarchy of the gold evidence
Figure 9: Sample Distribution of Multihiertt partitioned by the percentage of empty cells in the supporting table
A.3 Categorization of LLM prompting errors in numerical tabular reasoning
Basel III Standardized Transitional Basel III Advanced Transitional (in millions) Dec 31, 2017 Dec 31, 2016 Dec 31, 2017 Dec 31, 2016 Regulatory capital CET1 capital \$184,375 \$179,319 \$184,375 \$179,319 Tier 1 capital(a) \$184,375 \$179,341 \$184,375 \$179,341 Total capital \$195,839 \$191,662 \$189,419 \$184,637 Assets Capital ratios(c) CET1 13.8% 13.7% (e) 15.0% Tier 1(a) 13.8% 13.7% (e) 15.0% Total 14.7% 14.6% (e) 15.4% Tier 1 leverage(d) 8.7% 8.6% 8.7% Question: What is the sum of CET1 capital, Tier 1 capital, and Total capital in 2017? (in million) Response #: Domain Knowledge: In banking, CET1 capital, Tier 1 capital, and Total capital are used to measure a bank’s financial health and stability. CET1 (Common Equity Tier 1) capital consists mainly of common shares and retained earnings. Tier 1 capital is the core capital mainly comprising shareholders’ equity and disclosed reserves. Total capital includes both Tier 1 capital and Tier 2 capital, which consists of hybrid capital instruments and subordinated debt. Gold Evidences: Table shows JPMorgan Chase Bank, N.A. Basel III Standardized Transitional CET1 capital on Dec 31, 2017 as \$184,375 million. Table shows JPMorgan Chase Bank, N.A. Basel III Standardized Transitional Tier 1 capital on Dec 31, 2017 as \$184,375 million. Table shows JPMorgan Chase Bank, N.A. Basel III Standardized Transitional Total capital on Dec 31, 2017 as \$195,839 million. Steps: 1. Add CET1 capital and Tier 1 capital. Response 1: \$184,375 + \$184,375 = \$368,750 2. Add #1 and Total capital. Response 2: #1 + \$195,839 = \$564,589 The final answer is \$564,589 million. Ground-Truth Steps: \$184,375 + \$184,375 + \$195,839 + \$184,375 + \$184,375 + \$189,419 = \$1,122,758 Reason: The model missed adding JPMorgan Chase Bank, N.A. Basel III Advanced Transitional CET1 capital (\$184,375), Tier 1 (\$184,375), and Total Capital (\$189,419) in 2017.
Regulatory capital, assets, and risk-based capital ratios for JPMorgan Chase and its significant IDI subsidiaries under Basel III Standardized Transitional and Basel III Advanced Transitional at December 31, 2017, and 2016.
Figure 10: Error due to missing evidences
Year Life/Fin. Gen. Rtd. Fin. Asset Total Ins. Ins. Services Services Mngmt 2006 Fixed Maturities: Bonds available for sale \$67,994 \$288,018 \$1,357 \$29,500 \$386,869 Bonds held to maturity \$21,437 \$21,437 Bond trading securities \$1 \$10,835 \$10,836 Equity Securities: Common stocks \$4,245 \$8,705 \$226 \$80 \$13,256 Preferred stocks \$1,884 \$650 \$5 \$2,539 Other Assets: Mortgage and other loans \$17 \$21,043 \$2,398 \$4,884 \$76 \$28,418 Flight equipment \$39,875 \$39,875 Total Invested Assets \$116,059 \$428,825 \$185,404 \$68,280 \$3,373 \$801,941 Question: What is the proportion of Asset Management to the total Mortgage and other loans receivable, net of allowance in 2006? Response #: Domain Knowledge: Proportion refers to the relationship between two quantities, where one quantity is considered in relation to the other. To find the proportion of Asset Management to the total Mortgage and other loans receivable, net of allowance in 2006, we need to divide the value of Asset Management by the total value of Mortgage and other loans receivable, net of allowance in 2006. Gold Evidences: Table shows Asset Management of Total invested assets in 2006 is \$68,280. Table shows Mortgage and other loans receivable, net of allowance (in millions) in 2006 is \$28,418. Steps: 1. Divide Asset Management by the total Mortgage and other loans receivable, net of allowance in 2006. Response 1: $\frac{68280}{28418}=2.4024$ The final answer is 2.4024. Ground-truth steps: $\frac{4884.0}{28418.0}=0.172$ Reason: Asset Management in 2006 is taken as \$68,280.0 by the model instead of \$4,884.0 due to error in the extraction of the ground-truth.
Figure 11: Error due to wrong evidences
2013 2014 2015 2016 2017 Masco \$138.48 \$155.26 \$200.79 \$227.08 \$318.46 S&P 500 Index \$132.04 \$149.89 \$151.94 \$169.82 \$206.49 S&P Industrials Index \$140.18 \$153.73 \$149.83 \$177.65 \$214.55 S&P Consumer Durables & Apparel Index \$135.84 \$148.31 \$147.23 \$138.82 \$164.39 Question: What was the difference in percentage cumulative total shareholder return on Masco common stock versus the S&P 500 Index for the five-year period ended 2017? Response #: Domain Knowledge: Cumulative total shareholder return is a measure of the overall performance of a stock over a given period of time. It includes the change in the stock’s price as well as any dividends paid to shareholders during that time. Gold Evidences: The cumulative total shareholder return for Masco in 2017 is \$318.46. The cumulative total shareholder return for the S&P 500 Index in 2017 is \$206.49. Steps: 1. Divide the cumulative total shareholder return for Masco in 2017 (318.46) by the cumulative total shareholder return for the S&P 500 index in 2017 (206.49) Response 1: 318.46 / 206.49 = 1.5422 2. Subtract 1 from the quotient obtained in #1 Response 2: 1.5422 - 1 = 0.5422 3. Multiply the value obtained in #2 with 100 to find the percentage difference Response 3: 0.5422 * 100 = 54.22 The final answer is 54.22%. Ground-truth steps: [((318.46 - 100.0)/100.$0$ ((206.49 - 100.0)/100.0)*100.0] = 111.97 Reason: Wrong formula for return on investment
Figure 12: Example of an error caused due to insufficient domain knowledge
Accr. Yield Carrying Amt. Allowance (Beginning) (Beginning) (Beginning) Beginning Balance \$27 \$920 \$95 Purchases \$1 \$130 \$0 Disposals/Payments Received \$-11 \$-594 \$0 Accretion \$-44 \$44 \$0 Builds (Reductions) to the Allowance \$0 \$0 \$-18 Increase to Expected Cash Flows \$-2 \$19 \$0 FX/Other \$17 \$-50 \$0 Balance at December 31, 2010 \$116 \$469 \$77 Question: What percentage did this total represent in relation to the carrying amount? Response #: Domain Knowledge: To find the percentage of a value (A) in relation to another value (B), divide the first value (A) by the second value (B) and multiply it by 100. To find the combined total, add the two values. Gold Evidences: The carrying amount of the company’s purchased distressed loan portfolio at December 31, 2010, was \$392 million. The net allowance was \$77 million as of December 31, 2010. Steps: 1. Find the combined total of net allowance and the carrying amount. Response 1: 392 + 77 = 469 2. Calculate the percentage of net allowance in relation to the carrying amount. Response 2: $\frac{77}{392}\times 100=19.64$ 3. Calculate the percentage of the combined total in relation to the carrying amount. Response 3: $\left(\frac{469}{392}\right)\times 100=119.64$ The final answer is 119.64%. Ground-truth steps: [(77/(392 + 77))*100] = 16.42% Reason: By the phrase in relation to the question expected a ratio of $quantity_{1}/quantity_{2}$ but the model returned $quantity_{2}/quantity_{1}$
Figure 13: Example of an error caused due to question misinterpretation
2001 2000 Gain on Sales of Depreciable Properties \$45428 \$52067 Gain on Land Sales \$5080 \$9165 Impairment Adjustment \$-4800 (4800) \$-540 (540) Total \$45708 \$60692 Question: What is the percent change in general and administrative expense from 2000 to 2001? Response #: Domain Knowledge: To find the percent change between two values, we need to find the difference between the new and old values, divide the difference by the old value, and multiply the result by 100. Gold Evidences: General and administrative expense decreased from \$21.1 million in 2000 to \$15.6 million for the year ended December 31, 2001. Steps: 1. Calculate the change in general and administrative expense by subtracting the 2001 value (\$15.6 million) from the 2000 value (\$21.1 million). Response 1: 21.1 - 15.6 = 5.5 2. Divide the difference obtained in Step 1 by the general and administrative expense in 2000 (\$21.1 million). Response 2: $\frac{5.5}{21.1}=0.2607$ 3. Multiply the result obtained in Step 2 by 100 to get the percent change. Response 3: $0.2607\times 100=26.07$ The final answer is 26.07%. Ground-truth steps: [((15.6 - 21.1)/21.31)*100] = -26.07% Reason: Incorrect Instantiation of variables resulting in a value with a flipped sign
Figure 14: Example of an error caused due to incorrect instantiation of variables
2013 2014 2015 2016 2017 CME Group Inc. \$164.01 \$194.06 \$208.95 \$279.85 \$370.32 S&P 500 \$132.39 \$150.51 \$152.59 \$170.84 \$208.14 Peer Group \$176.61 \$187.48 \$219.99 \$249.31 \$323.23 Question: What is the annualized return for s&p 500 from 2012 to 2017? Response #: Domain Knowledge: The annualized return refers to the average rate of return over a given period for an investment, which has been adjusted for compounding. To find the annualized return for the S&P 500 from 2012 to 2017, we need to use the formula: Annualized $\text{Annualized return}=\left(\left(\frac{\text{Ending value}}{\text{Startingvalue}}\right)^{\frac{1}{\text{Number of years}}}-1\right)$ Gold Evidences: Investment was \$100 on December 31, 2012;Value of S&P 500 investment on December 31, 2017 is \$208.14. Steps: 1. Divide the ending value (208.14) by the starting value (100) Response 1: $\frac{208.14}{100}=2.0814$ 2. Calculate the number of years elapsed (2017 - 2012) Response 2: 2017 - 2012 = 5 3. Raise the value obtained in #1 to the power of (1/number of years) Response 3: $(2.0814)^{\frac{1}{5}}=1.1606$ 4. Subtract 1 from the value obtained in #3 Response 4: 1.1606 - 1 = 0.1606 5. Multiply the value obtained in #4 by 100 to get the annualized return in percentage Response 5: 0.1606 * 100 = 16.06 The final answer is 16.06%. Ground-truth steps: [$(208.14/100)^{(}1/(2017-2012))$ 1] = 0.1577 * 100 = 15.77 Reason: Inaccurate calculation
Figure 15: Example of an error caused due to imprecise calculations
Figure 16: This illustration demonstrates how the "depth of hierarchy" is determined for each supporting piece of evidence. The depth of hierarchy corresponds to the number of levels or tiers of information structure within the table. The top level (1) encompasses the table itself, which contains information about the results of operations in a firm’s Advice & Wealth Management segment. The second level (2) includes columns specifying the years (2009, 2008, Change), creating a substructure within the table. The third level (3) involves the column "(in millions, except percentages)," and the fourth level (4) encompasses the rows under categories like "Revenues," "Distribution fees," "Net investment income," and so on. To locate specific values like "\$1733" and "\$1912", a retriever module needs to navigate through these four levels.
Figure 17: A Template for our proposed prompting strategy, EEDP with 1-shot demonstration.